### v1

# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

Abstract.
The ability to research and synthesize knowledge is central to human expertise and progress. An emerging class of systems promises these exciting capabilities through generative research synthesis, performing retrieval over the live web and synthesizing many discovered sources into long-form, cited summaries. However, evaluating such systems remains an open challenge: existing question-answering benchmarks focus on short-form factual responses, while expert-curated datasets risk staleness and data contamination. Both fail to capture the complexity and evolving nature of real research synthesis tasks. In this work, we introduce DeepScholar-bench, a live benchmark and holistic, automated evaluation framework designed to evaluate generative research synthesis. DeepScholar-bench draws queries from recent, high-quality ArXiv papers and focuses on a real research synthesis task: generating the related work sections of a paper by retrieving, synthesizing, and citing prior research. We develop an automated evaluation framework that holistically assesses performance across three key dimensions, knowledge synthesis, retrieval quality, and verifiability, using metrics that show strong agreement with expert human judgments. We also develop DeepScholar-base, a reference pipeline for generative research synthesis, implemented efficiently using the LOTUS API. Using the DeepScholar-bench framework, we perform a systematic evaluation of prior open-source systems, search AI’s with open-source and strong proprietary models, OpenAI’s DeepResearch, and DeepScholar-base. We find that DeepScholar-base establishes a strong baseline, attaining competitive or higher performance than prior open-source systems, search AI’s and OpenAI’s DeepResearch. We also find that DeepScholar-bench remains far from saturated, with no system exceeding a score of across all metrics. These results underscore both the difficulty and the importance of DeepScholar-bench as a foundation for progress toward AI systems capable of generative research synthesis. We make our benchmark code and data available at https://github.com/guestrin-lab/deepscholar-bench.
1. Introduction
A crucial underpinning of human knowledge and innovation is the ability of human experts to research and synthesize known facts and new findings to support others in comprehending, verifying and building upon existing work. Recently, new systems designed for generative research synthesis offer exciting capabilities and promises to automate challenging research synthesis tasks, which often require laborious hours of literature search, reading, and writing from human experts. These systems include a wide array of both commercial DeepResearch offerings, including ones from OpenAI (noa, [n.d.]g), Gemini (noa, [n.d.]c), Anthropic (noa, [n.d.]b), Grok (noa, [n.d.]e), and Perplexity (noa, [n.d.]h), as well as open-source methods, such as STORM (Shao et al., 2024), DeepResearcher (Zheng et al., 2025), and OpenScholar (Asai et al., 2024). These systems push the frontier of AI capabilities, demonstrating promising performance on existing factuality and question-answering benchmarks (Wei et al., 2024; Krishna et al., 2025; Mialon et al., 2023; Wei et al., 2025).
Yet, as this new class of systems emerges, a key question remains: how should we benchmark and evaluate generative research synthesis? The progress of these systems requires benchmarks that carefully evaluate their critical capabilities. Specifically, these systems provide three core functions: (1) retrieval, typically from a large, complex, and constantly-evolving corpus, such as the live web, to collect key information (2) knowledge synthesis, to generate coherent long-form answers that surface key facts, integrating general knowledge and findings from many retrieved sources, and (3) verifiability, providing citations that allow readers to trace each stated claim in the synthesized answer to a reputable source from the retrieved set. The ideal benchmark must holistically evaluate across all three of these dimensions, while providing a realistic and challenging research synthesis task, which the community can reliably and repeatedly benchmark on over time.
Unfortunately, existing benchmarks fall-short of these goals. Many prior works evaluate generative research synthesis systems using existing question answering benchmarks (Wei et al., 2025, 2024; Mialon et al., 2023; Krishna et al., 2025; Wu et al., 2025; Wadden et al., 2020; Jin et al., 2019; Yang et al., 2018; Joshi et al., 2017; Kwiatkowski et al., 2019; Ho et al., 2020; Trivedi et al., 2022; Lee et al., 2023), which do not reflect realistic research synthesis tasks and instead focus on questions with short-form, easily-verifiable answers, making them severely limited for this setting. These question-answering benchmarks do not capture the complexity of long-form answers synthesized from many sources, a key component of research synthesis. To address this limitation, several recent works (Asai et al., 2024; Zheng et al., 2024; noa, [n.d.]k) instead leverage expert-curated datasets with open-ended research questions and exemplar answers. Unfortunately, these benchmarks quickly become stale and outdated as new information emerges. Furthermore, these datasets risk data contamination as new models are trained on snapshots of the web, including public datasets. The prohibitive expense of curating, maintaining, and updating expert-curated benchmarks further limits their utility towards realistic, scalable evaluation.
To address this gap, we introduce DeepScholar-Bench, a live benchmark dataset and holistic evaluation framework tailored to tracking progress on generative research synthesis, as well as DeepScholar-Base, a reference method, which we hope will further future research. Our benchmark provides a live dataset with real research synthesis tasks, as well as a holistic, automated evaluation strategy. We design our benchmark dataset based on a fundamental research synthesis task reflected in thousands of research papers each year, where peer-reviewed manuscripts must concisely summarize existing work within a subfield, detailing and citing key prior works, before discussing new contributions. Specifically, the DeepScholar-bench task requires a system to generate a related work section for a given academic paper by retrieving sources from the Web and summarizing key prior work within a field of study. This task allows us to assess the key capabilities of research synthesis while ensuring that our evaluation reflects real, high-quality research. To source realistic and challenging queries, we develop an automated data pipeline that curates high-quality, recent ArXiv papers across diverse scientific domains. We plan to continuously re-run our data pipeline to provide an evolving dataset of live queries.
To assess performance on DeepScholar-bench, we develop an automated evaluation framework, which holistically measures the performance of system answers across the three key dimensions: knowledge synthesis, retrieval quality and verifiability. Evaluating the accuracy of our long-form synthesis task is inherently difficult since each query admits many possible answers, lacking a straightforward notion of correctness. Moreover, developing an automated evaluation requires reliable metrics that exhibit high agreement with expert human annotators, another significant challenge. To address these challenges, our holistic evaluation assesses each system response across seven key metrics (Table 1), which permit many possible correct answers, often leveraging human-written exemplars from our dataset. Specifically, on the knowledge synthesis dimension, we evaluate a generated answer’s Organization, using pairwise comparisons to human exemplars, and Nugget Coverage, assessing the efficiency of the generated response in capturing key information and facts. To assess retrieval quality, we measure the Relevance Rate of retrieved sources, the Document Importance of sources, according to citation counts of each reference, and Reference Coverage, by assessing the generated report’s coverage of notable important references recovered from the human-written exemplar. To assess the Verifiability of each report, we measure its Citation Precision, whether each citation supports the given claim, and Citation Coverage, measuring whether each claim is fully supported by the cited sources. Our human agreement study validates that these automated metrics are effective, demonstrating strong agreement between our model-based judges and expert human annotators.
| Metric | Description |
| Knowledge Synthesis | |
| Organization | assesses organization and coherence of system answer |
| Nugget Coverage | assesses the answer’s coverage of essential facts |
| Retrieval Quality | |
| Relevance Rate | measures avg. relevance among all referenced sources |
| Document Importance | measures how notable referenced sources are, using citation counts |
| Reference Coverage | assesses the referenced set’s coverage of key, important references |
| Verifiability | |
| Citation Precision | measures percent of cited sources that support their accompanying claim |
| Claim Coverage | measures percent of claims that are fully supported by cited sources |
Using the DeepScholar-bench framework, we systematically evaluate the performance of existing systems, including open-source research synthesis systems, search AI’s with strong proprietary models, OpenAI DeepResearch, and DeepScholar-base. Figure 1 demonstrates that all of these existing methods exhibit significant opportunity for improvement, with no baseline exceeding a score of across all metrics. Furthermore, on several key metrics, including Nugget Coverage, Reference Coverage and Document Importance, each evaluated method’s performance remains well below , reflecting the inherent difficulty of the DeepScholar-bench task, which requires systems to navigate the live web, reasoning about the relevance and importance of documents as well as surfacing key facts into a concise final answer. Notably, OpenAI’s DeepResearch offers strong performance relative to other baselines, outperforming many prior methods on knowledge synthesis and retrieval quality, with scores of on Nugget Coverage, on Reference Coverage and on Document Importance; however, it struggles to provide strong verifiability relative to many other methods. We also find that DeepScholar-base, a relatively simple reference pipeline implemented efficiently using the LOTUS API (noa, 2025b; Patel et al., 2025), consistently improves upon performance of prior open-source systems and search AI’s, as shown in Figure 1, as well as achieving competitive performance and up to higher verifiability compared to OpenAI’s DeepResearch, as shown in Figure 1. Nevertheless, DeepScholar-bench remains far from saturated, representing exciting opportunities for further work. We hope that our benchmark framework and reference pipeline support the progress of new systems, and we believe that resolving DeepScholar-bench represents a critical milestone towards more capable AI systems.
Overall, our main contributions are the following:
We propose DeepScholar-bench, a live benchmark dataset with real research synthesis tasks and an automated evaluation that holistically assesses performance across three key dimensions: knowledge synthesis, retrieval quality, and verifiability. Our analysis demonstrates that our DeepScholar-bench evaluation metrics are robust, exhibiting strong human agreement scores.
We develop DeepScholar-base, a reference pipeline for generative research synthesis, providing a strong baseline with competitive or higher performance compared to prior open-source systems, search AI’s, and OpenAI’s DeepResearch.
We perform a systematic evaluation of existing baselines on DeepScholar-bench, including open-source systems, search AI’s, OpenAI’s DeepResearch, and DeepScholar-base. We find that DeepScholar-bench is far from saturated, demonstrating significant opportunity for improvement, with no baseline achieving a score greater than across all metrics.
2. The DeepScholar Dataset
We study the task of generating a related works section of an academic paper, a fundamental research synthesis task. We choose this task for two key reasons. First, this task is a real research task performed by academic experts, allowing our benchmark to reflect realistic, difficult and useful queries. Second, the online availability of diverse, high-quality academic papers allows us to develop an automated dataset construction pipeline that we can continuously run to obtain new queries over time.
We construct our dataset by scraping papers from ArXiv (noa, [n.d.]a), which continuously posts thousands of new pre-print papers across a wide array of scientific domains each week. We formalize our dataset task as follows: given a description, of a paper, the goal is to retrieve a set of relevant sources, , and generate a related works sections, , for the paper by synthesizing and citing the retrieved documents. We briefly overview our automated data collection framework (Section 2.1) and describe the dataset instantiation (Section 2.2) used in our evaluation (Section 5).
2.1. Automated Data Collection Framework
Our automated data collection framework aims to achieve the following design goals:
Inclusion of diverse paper topics across a wide variety of research domains.
Focus on recent research papers, both to provide realistic, timely benchmark queries and to control data contamination when benchmarking models trained on snapshots of the web.
Control for quality of the scraped ArXiv papers and extracted data
Figure 2 provides an overview of our dataset pipeline, which collects and extracts metadata about each paper (e.g., the title, abstract, and ArXiv link), the paper’s related works section, and information on each citation from the paper’s related works section. Our data collection pipeline extracts this information by scraping and selectively filtering ArXiv papers according to a number of configured settings.
Specifically, the pipeline loads papers from a list of configured ArXiv domain categories (e.g., cs.ML) and filters paper according to the configured publication-date range. To avoid possible data contamination arising from multiple ArXiv versions, some of which may have been released prior to the configured publication-date range, we exclusively include v1 ArXiv papers. To control for paper quality, our pipeline optionally provides a configuration setting which filter’s out papers which are not listed as ”accepted” or ”published” at a conference within the paper’s comment metadata, which often lists updates to the paper’s status. We also disclude papers that do not have an explicit ”Related Works” section and .bib file, containing well-formatted bibliography entries. For each paper, we then extract the Related Works section, from both the LaTex files, and PDF file, if available. We clean the extracted LaTex section to remove figures, sub-figures, labels and comments. We also extract all citations found in the related work section from the LaTex .bib file. For each citation we use the ArXiv API and the OpenAlex API to recover more detailed information, such as abstracts, authors, and links for ArXiv and non-ArXiv references respectively.
2.2. Dataset Description and Statistics
We now briefly summarize the dataset we use in our evaluation in Section 5 which represents an instantiation of our automated data collection pipeline. Our datasets take ArXiv papers with a publication date between April and June, following the April 5th, 2025 release date of the Llama-4 models (noa, 2025a), the main open-source model we benchmark in our evaluation. Our dataset consist of papers scraped from a diverse set of 18 disticnt ArXiv domains, including, cs.AI, cs.CV, cs.DB, cs.LG, cs.AR, cs.CG, cs.DC, cs.DS, cs.IR. To control for quality, we filter out papers not accepted at a conference, and we additionally exclude papers with related works sections longer than 1,000 words. Our final dataset instantiation includes 63 ArXiv papers, each providing a single query and expert-written exemplar for our benchmark. We make our scripts available to allow others to configure different datasets, and we plan to update our dataset to provide a continual evaluation with recent queries. Our experiments leverage the abstract of each paper as the paper’s description , provided to each baseline system as context within the query. We analyze the human-written exemplars from our dataset, and we find that, on average, each related work section contains unique references, and we find over of all cited references on ArXiv.
3. The DeepScholar Evaluation Framework
Evaluating research synthesis systems is challenging due to the complexity of the task and the variability of possible correct answers. Research synthesis systems generate complex, long-form reports, which are difficult to evaluate and lack a notion of ”ground truth” correctness. The task we consider departs significantly from traditional question answering and RAG-based evaluations (Joshi et al., 2017; Lee et al., 2023; Jin et al., 2019; Kwiatkowski et al., 2019; Trivedi et al., 2022; Yang et al., 2018; Ho et al., 2020), requiring a carefully designed and holistic evaluation framework. An exemplar research report must retrieve important relevant sources, synthesize an informative and well-organized answer, and provide appropriate references allowing readers to verify and re-trace facts. Our holistic evaluation framework thus analyzes three key dimensions, providing an automated, scalable approach for each: knowledge synthesis (Section 3.1), retrieval quality (Section 3.2), and verifiability (Section 3.3). While our experiments in Section 5 evaluate one specific task, our evaluation framework may extend to a wide range of research synthesis tasks (Shao et al., 2024; noa, [n.d.]k; Zheng et al., 2024; Asai et al., 2024), which exhibit similar properties and challenges. We provide an overview of our evaluation framework in this section and further details and analysis in Appendix Section 8.4.
3.1. Knowledge Synthesis
We evaluate both the information content surfaced in each synthesized report and the overall organization and coherence of the report.
Organization and Coherency. Our automatic evaluation adopts an LLM-as-a-judge approach to assess the organization and coherence of each system answer. We use preference-based pairwise comparisons, where the LLM-judge is presented with the details of the criteria to judge, the human-written exemplar, and a generated report and is asked to mark which is better. To avoid position bias (Li et al., 2025), we evaluate each pair twice, permuting their positions. This model-based evaluation provides scalability while also serving as a strong surrogate for human preferences (Rahmani et al., 2024b; Li et al., 2024, 2025; Arabzadeh and Clarke, 2025), which we validate in our experiments in Section 5. We report the win-rate of each system using the prompt shown in Box 8.4.3 in the Appendix.
Nugget Coverage. To assess the quality of the information content presented by a generated report, we use a nugget-based evaluation. An information nugget (Pradeep et al., 2025; Upadhyay et al., 2024b; Faggioli et al., 2023; Rahmani et al., 2024a; Upadhyay et al., 2024a) is an essential fact or components relevant for an answer. The process of nuggetization involves decomposing information-dense text into essential components, which aid in evaluation. For our task, we generate nuggets from the human-written exemplar related-work section for each query, following the automated, LLM-based methodology of Pradeep et al. (Pradeep et al., 2025). In section 5 we validate that the model-based approach has strong agreement with expert annotations when labeling nuggets. For, each report, we compute the nugget coverage score, the fraction of nuggets that are present in each system answer.
3.2. Retrieval Quality
While traditional information retrieval (IR) evaluations typically leverage gold labels for document relevance scores and a controlled corpus (Thakur et al., 2021; noa, [n.d.]l; Nanni et al., 2017), the research synthesis task we study in this work differs substantially. An expert-written report section and its sources provide one reasonable retrieved set, but there may be many possible alternative sets that are likewise high-quality. Moreover, live web search is a key component of research synthesis tasks and obtaining gold relevance labels over this corpus is prohibitively expensive. To address these challenges, our evaluation measures three components of the retrieved set: the relevance rate, reference coverage of key sources, and document importance.
Relevance Rate. We asses the relevance of each retrieved document, following the Cranfield model (Voorhees, 2009), which is standard in IR evaluations and considers relevance of individual documents given a query, independent of other documents. Due to the significant cost of obtaining human-annotated relevance judgments, recent works (Upadhyay et al., 2024b; Faggioli et al., 2023; Rahmani et al., 2024b; Thomas et al., 2024; Asai et al., 2024) study leveraging an LLM-as-a-judge for relevance judgment task, demonstrating their effectiveness on traditional IR datasets. Building on these works, we adopt an LLM-based approach for assigning graded relevance scores from 0 to 2 to each generated research report using the prompt in Box 8.4.3 in Appendix. For each retrieved set, , for a given query, we compute the average document relevance over the set, following the below formula:
where is the graded relevance score of source . We validate the agreement between LLMs and human annotators for this task in our experiments in Section 5.
Reference Coverage. We introduce a metric to measure the reference coverage of the retrieved set for each report. A key challenge in measuring this value is in defining a set of important references for each generated report, that a good research report should cite. To build this set, we take all references from the high-quality, human-written exemplar and label each as either ”important” or ”not-important”, considering a ”not-important” reference as one that could be omitted or substituted by a different reference. We find that a LLM-based judge is effective in assigning these labels. For a given report, we then compute its reference coverage by taking the ratio of the number of important references in the system-generated report to the number of important references in the human-written exemplar, following the below formula, where is the set of ”important” references from the human-written exemplar:
Document Importance. While the above relevance and coverage metrics assess topical matches between the retrieved set and the user query, an ideal research synthesis system must also retrieve notable and important sources. Exemplar human-written reports typically contain ample references of primary-sources and highly-cited academic publications. While the ideal notion of document importance depends on the particular task and user, in our task, we adopt the following metric by considering the number of citations that each reference retrieved by the system has. We consider the median number of citations per reference over the set of all retrieved sources, , provided by a given baseline. We compute document importance as the ratio of this value for the given baseline compared to the median citations per reference over the set of sources, provided by the human-written exemplars, following the formula below:
where is the number of citations for source . We set an upper-bound of one, although in practice, we find this ratio to remains far below one for all measured generative research synthesis systems.
3.3. Verifiability
To evaluate the verifiability of the generated report given the retrieved set, we calculate the citation precision and claim coverage based on the definitions provided by prior work (Gao et al., 2023; Worledge et al., 2024; Liu et al., 2023).
Citation Precision. Specifically, we measure sentence-level precision, where a citation is considered precise if the referenced source supports at least one claim made in the accompanying sentence. For a full report, citation precision is computed by averaging the precision of each citation in the report.
Claim Coverage. Claim coverage assigns a sentence-level score, assigning a value of one if the cited sources accompanying the sentence supports all claims made in the sentence. We make two adaptations to the original definition posed in prior work (Gao et al., 2023; Worledge et al., 2024; Liu et al., 2023) for our long-form synthesis task. First, we relax the original claim coverage definition to consider a sliding window of sentences with supporting references, assigning a coverage value of 1 to each sentence that is either fully supported by the sources cited within the sentence or any cited source that is in a window of preceding or following sentences. Additionally, since our task query provides context describing a paper, we consider this context as an implicitly cited reference for each sentence and compute claim coverage for each sentence by considering the explicitly cited sources, and the context provided by the query. We compute the citation coverage for the full report by averaging the value computed for each sentence. Our model-based evaluation uses an LLM judge to assess each entailment relation, following prior work (Gao et al., 2023; Liu et al., 2023). We provide further details and the prompt in the Appendix Section 8.4 and Box 8.4.3 respectively.
4. DeepScholar-base
We introduce DeepScholar-base, an open-source reference pipeline designed to perform generative research synthesis. Figure 3 provides an overview of our method. Given a user’s query, DeepScholar-base iteratively generates web-search queries, summarizes the search results in each round before generating a new query. The system then post-processes the search results leveraging a series of semantic operators (Patel et al., 2025), which we implement efficiently using the LOTUS API (noa, 2025b). This includes a semantic filtering step, which leverages an LLM to filter out irrelevant source documents, followed by a semantic top-k which performs an LLM-based ranking over the documents based on their relevance to the user query. Finally, we perform a semantic aggregation over the final source documents to generate the final report. We provide further details of each step of our reference pipeline in Appendix Section 8.3.
5. Experimental Results
In this section, we evaluate recent state-of-the-art generative research systems as well as DeepScholar-base on DeepScholar-bench. Overall, we find the following:
Existing baselines for generative research synthesis, including strong open-source LLM systems, search AI’s, and commercial systems, demonstrate significant room for improvement across all three key dimensions: knowledge synthesis, retrieval quality and verifiability, with no method exceeding a score of across all metrics.
DeepScholar-base provides a strong baseline for generative research synthesis, consistently improving upon the performance of prior open-source systems and search AI’s, as well as achieving competitive performance and up to higher verifiability compared to OpenAI’s DeepResearch.
Our automated evaluation approach is effective, demonstrating high agreement with over 200 human annotations.
Experimental Setup. For each benchmarked method, we control the retrieval corpus by allowing each system to access the Web only through the ArXiv API (noa, [n.d.]a). We additionally avoid possible information leakage during search by filtering out any search results that were published after the query paper’s publication date. We report results using GPT-4.1-2025-04-14 (noa, [n.d.]i) as the judge for Nugget Coverage, and a GPT-4o-2024-08-06 (noa, [n.d.]i) judge for Organization, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage. We report the Organization score as a win rate including ties, we report the strict all score for Nugget Coverage, and we report Claim Coverage with a window size of . For all Retrieval Quality metrics, we consider the retrieved set of each given report as the set of any valid ArXiv links found within the report. To measure Document Importance, we use the OpenAlex (noa, [n.d.]j) API to recover citation information. For each metric, we report an average over all reports.
5.1. Baselines
We briefly overview all of the baseline systems we evaluate, and we provide further details in the Appendix Section 8.2.
5.1.1. Open-source Research Systems
We evaluate three state-of-the-art open-source systems, DeepResearcher (Zheng et al., 2025), STORM (Shao et al., 2024) and OpenScholar (Asai et al., 2024). For each, we run these systems using the Llama-4-Scout-17B-16E-Instruct model (noa, 2025c), which we serve with 4 A100 GPUs using vLLM (Kwon et al., 2023).
DeepResearcher (Zheng et al., 2025) leverages trained agents to navigate, browse and synthesize information from the web. To train an agent, this work uses end-to-end reinforcement learning and trains Qwen2.5-7B-Instruct (Qwen et al., 2025). In our benchmarks, we evaluate DeepResearcher using both the released, trained model from the authors, and using Llama-4-Scout-17B-16E-Instruct model (noa, 2025c) as the core LLM. We report the better performing baseline of these two, which we find in our experiments to be the Llama-4-Scout-17B-16E-Instruct backbone.
STORM (Shao et al., 2024) studies the problem of how to apply LLMs to write grounded, organized long-form articles (e.g., Wikipedia articles) from scratch. The system involves a pre-writing stage that discovers diverse research perspectives on a topic by stimulating conversations between multiple agents and leveraging web documents.
OpenScholar (Asai et al., 2024) builds a specialized retrieval-augmented LLM system for literature synthesis and scientific queries. This method includes a trained retriever from the pre-indexed peS2o (Soldaini et al., 2024) corpus, consisting of 45 million open-access academic papers up until October 2024, as an initial retrieval source before using web search. In our experiments, we benchmark the system using this pre-indexed corpus and limit web search to the ArXiv API.
5.1.2. Search AI’s
We evaluate the following models: Llama-4-Scout-17B-16E-Instruct (noa, 2025c), GPT-4.1-2025-04-14 (noa, [n.d.]i), o3-2025-04-16 (noa, [n.d.]i), Claude-opus-4-20250514 (noa, [n.d.]f), and Gemini-2.5-pro (noa, [n.d.]d). We augment each with search capabilities to ArXiv (noa, [n.d.]a), and use the popular ODS framework (Alzubi et al., 2025) to allow the LLM to make tool calls to the search API.
5.1.3. Commercial Systems.
We focus our evalution of commercial generative research synthesis systems on OpenAI’s o3-deep-research (noa, [n.d.]i), which provides a public API allowing for our evaluation.
5.1.4. DeepScholar-base
Similar to our evaluation of search AI’s we evaluate DeepScholar-base with the following models: Llama-4-Scout-17B-16E-Instruct (noa, 2025c), GPT-4.1-2025-04-14 (noa, [n.d.]i), o3-2025-04-16 (noa, [n.d.]i), Claude-opus-4-20250514 (noa, [n.d.]f), and Gemini-2.5-pro (noa, [n.d.]d). For each of these baselines, we also use the same or a weaker model, either Llama-4 or GPT-4.1, to perform semantic filtering and top-k operators. We limit the method to two round of search, each with at most 2 queries.
5.2. Main Results
Table 2 provides detailed summary of each method’s performance scores on all metrics across three key dimensions, knowledge synthesis, retrieval quality and verifiability. We also provide metadata statistics characterizing the generated reports of each benchmarked method in Table 5, as well as statistics related to our evaluation metrics in Table 6. Overall, our evaluation demonstrates two key findings, which we discuss in detail below: first, existing generative research synthesis systems demonstrate significant headway for improvement, and second, DeepScholar-base provides a strong baseline for generative research synthesis.
| Knowledge Synthesis | Retrieval Quality | Verifiability | |||||
| Org. | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | |
| Human-written Exemplars | |||||||
| Human-written Exemplars | .500 | 1.000 | .585 | 1.000 | 1.000 | .278111The verifiability metrics we use in our evaluation likely significantly under-estimate the actual verifiability of human writing. This is because the metrics we measure, Citation Precision and Claim Coverage, require us to verify entailment relations between claims, within the report, and snippets from the cited reference. For each LLM-based system, we are able to track the precise snippet and context from each cited source, which are directly fed as context to the LLM as it generates the report with citations. On the other hand, for the human-written exemplars, we lack gold labels pointing to the precise snippet of text that each reference refers to. Our measurements for the human-written exemplars instead rely on the title and abstract of each cited source when testing for entailment relations. | .205111The verifiability metrics we use in our evaluation likely significantly under-estimate the actual verifiability of human writing. This is because the metrics we measure, Citation Precision and Claim Coverage, require us to verify entailment relations between claims, within the report, and snippets from the cited reference. For each LLM-based system, we are able to track the precise snippet and context from each cited source, which are directly fed as context to the LLM as it generates the report with citations. On the other hand, for the human-written exemplars, we lack gold labels pointing to the precise snippet of text that each reference refers to. Our measurements for the human-written exemplars instead rely on the title and abstract of each cited source when testing for entailment relations. |
| Open Source Research Systems | |||||||
| DeepResearcher (Llama-4) | .206 | .230 | .385 | .047 | .008 | .312 | .396 |
| STORM (Llama-4) | .119 | .183 | .218 | .003 | .006 | .238 | .586 |
| OpenScholar (Llama-4) | .309 | .278 | .017 | .008 | .013 | .010 | .138 |
| Search AI’s | |||||||
| Search AI (Llama-4-Scout) | .151 | .193 | .445 | .060 | .009 | .316 | .368 |
| Search AI (GPT-4.1) | .556 | .265 | .490 | .050 | .009 | .498 | .470 |
| Search AI (o3) | .849 | .348 | .610 | .165 | .026 | .425 | .495 |
| Search AI (Claude) | .698 | .307 | .583 | .131 | .008 | .701 | .760 |
| Search AI (Gemini) | .706 | .277 | .583 | .061 | .010 | .415 | .398 |
| Commercial Systems | |||||||
| OpenAI DeepResearch | .857 | .392 | .629 | .187 | .124 | .399 | .138 |
| DeepScholar Baseline | |||||||
| DeepScholar-base (Llama-4) | .254 | .262 | .421 | .103 | .008 | .648 | .826 |
| DeepScholar-base (GPT-4.1) | .825 | .407 | .608 | .162 | .007 | .652 | .636 |
| DeepScholar-base (GPT-4.1, o3) | .857 | .405 | .659 | .162 | .008 | .617 | .614 |
| DeepScholar-base (GPT-4.1, Claude) | .786 | .370 | .586 | .167 | .007 | .936 | .817 |
| DeepScholar-base (GPT-4.1, Gemini) | .762 | .332 | .602 | .181 | .006 | .851 | .865 |
5.2.1. Generative Research Synthesis Systems Demonstrate Large Room for Improvement.
From Table 2, we see that no method is able to achieve a score greater than across all metrics. Moreover, on several key metrics, including nugget coverage, reference coverage and document importance, each evaluated method’s performance remains well below . This reflects the inherent difficulty of the generative research task provided by DeepScholar-bench, which requires systems to navigate the live web, reasoning about relevance and importance of documents to perform retrieve sources and then surface key information into a coherent report that answers the query.
We now analyze each evaluated dimension, comparing performance of the open-source research systems, search AI’s and commercial systems to the human-written exemplars. On knowledge synthesis, we see that OpenAI DeepResearch offers the best performance compared to all other prior methods on both Organization, with a score of , and Nugget Coverage, with a score of . OpenAI DeepResearch, as well as the o3, Claude and Gemini search AI’s achieve relatively high Organization scores compared to human-written exemplars. However, on Nugget Coverage all prior methods scores below . This demonstrates that while existing systems, especially those using state-of-the-art models, can generate well-organized and coherent summaries, they still struggle to extract and surface key facts, a crucial capability for synthesis tasks.
Turning our attention to the retrieval quality performance of prior methods, we once again find significant room for improvement. Once again, OpenAI DeepResearch offers the strongest performance among the other benchmarked prior methods on Relevance Rate, Reference Coverage and Document Importance, but still far from saturates performance. While it’s Relevance Rate shows strong performance, exceeding that of the human exemplars with a score of , it’s Reference Coverage and Document Importance scores remain exceedingly low: and respectively. This demonstrates that while state-of-the-art generative research synthesis systems are capable of retrieving relevant sources, they still struggle to find a comprehensive set of notable sources and fall short compared to the ability of human experts.
Lastly, we analyze the verifiability performance of prior methods, we see that OpenAI DeepResearch is outperformed on both Citation Precision and Claim Coverage by the search AI’s with GPT4.1, o3, Claude and Gemini models. The Claude search AI offers the highest Citation Precision, a score of and Claim Coverage, a score of . Meanwhile, OpenAI’s DeepResearch as well as the all other prior methods are unable to achieve a Citation Precision score beyond and a Claim Coverage score beyond . We also note that the human-written exemplars appear to exhibit rather low Citation Precision and Claim Coverage scores, however these scores are not comparable to the metric measured for the LLM-based systems since our method for measuring verifiability metrics likely under-estimate the actual verifiability of human writing11footnotemark: 1. Overall, we see that prior LLM-based systems exhibit significant headroom for improvement.
| Knowledge Synthesis | Retrieval Quality | Verifiability | |||||
| Org | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | |
| DeepScholar-base (GPT-4.1, Claude) | |||||||
| arxiv.org Retrieval | .786 | .370 | .586 | .167 | .007 | .936 | .817 |
| parallel.ai Retrieval | .865 | .444 | .675 | .160 | .017 | .846 | .781 |
| taviliy.com Retrieval | .929 | .327 | .550 | .070 | .015 | .711 | .578 |
| Oracle Retrieval (arxiv.org) | .782 | .487 | .686 | 1.000 | 1.000 | .955 | .899 |
| Oracle Retrieval (All) | .778 | .528 | .680 | 1.000 | .822 | .941 | .828 |
| DeepScholar-base (Llama-4) | |||||||
| arxiv.org Retrieval | .254 | .262 | .421 | .103 | .008 | .648 | .826 |
| parallel.ai Retrieval | .246 | .265 | .559 | .114 | .015 | .223 | .543 |
| taviliy.com Retrieval | .111 | .229 | .532 | .030 | .016 | .442 | .676 |
| Oracle Retrieval (arxiv.org) | .202 | .316 | .681 | 1.000 | 1.000 | .658 | .868 |
| Oracle Retrieval (All) | .198 | .350 | .693 | 1.000 | .822 | .796 | .890 |
| Evaluation Metric | LLM-Classified Labels | Human-Agreement Score with LLM |
| Organization | Pairwise Comparison (Lose / Tie / Win) | |
| Nugget Coverage | Nugget Importance (Vital / Non-vital) | |
| Retrieval Relevance Score | Graded Relevance (0/1/2) | |
| Reference Coverage | Reference Importance (Not Imp./ Imp.) |
5.2.2. DeepScholar-base Provides a Strong Baseline for Generative Research Synthesis.
We compare the performance of DeepScholar-base to the commercial OpenAI DeepResearch system, search AI’s and open-source research systems, finding that DeepScholar-base provides competitive performance against each group of prior methods, offering a strong baseline for generative research synthesis.
First, we see that DeepScholar-base (GPT-4.1, o3) is competitive with OpenAI DeepResearch, achieving a similar or higher Organization, Nugget Coverage, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage scores. Notably, DeepScholar-base offers significantly higher verifiability than OpenAI DeepResearch, with higher Citation Precision and higher Claim Coverage. However, DeepScholar-base’s document importance scores still remain especially low compared to OpenAI DeepResearch, representing significant room for improvement.
Next, we compare the performance of DeepScholar-base to the search-AI’s, finding that for each evaluated model, DeepScholar-base consistently offers improved performance compared to the corresponding search AI. Specifically, averaged across all 5 baselines with different models for the Search AI’s and DeepScholar-base method, DeepScholar-base offers higher Organization, higher Nugget Coverage, higher Relevance Rate, higher Reference Coverage higher Citation Precision, higher Citation Recall.
Lastly, we compare DeepScholar-base (Llama-4) to the open-source research systems, all run with the Llama-4 model as well. We see that the prior open-source research systems exhibit trade-offs among the Knowledge Synthesis, Retrieval Quality and Verifiability dimensions. Specifically, OpenScholar achieves the highest Knowledge Synthesis scores, on both Organization and Nugget Coverage, DeepResearcher achieves the highest Relevance Rate and Reference Coverage on Retrieval Quality, with all systems attaining only very low Document Importance scores, and on Verifiability, DeepResearcher offers the highest Citation-Precision while STORM offers the highest Claim Coverage. In comparison, DeepScholar-base offers strong performance across each dimension. Specifically, Compared to the best-performing prior open-source methods for each metric, DeepScholar-base offers competitive Knowledge Synthesis performance, higher Relevance Rates and higher Reference Coverage for retrieval, and higher Citation Precision and higher Claim Coverage scores for verifiability.
Overall, the strong relative performance of DeepScholar-base likely reflects the efficiency of the data-processing semantic operators (Patel et al., 2025) that DeepScholar-base uses to perform LLM-based filtering, ranking and summarization of sources to generate its report. Notably, DeepScholar-base still demonstrates significant room for improvement and far from saturates DeepScholar-Bench, especially on key Knowledge Synthesis and Retrieval Quality metrics, including Nugget Coverage, Reference Coverage, Document Importance, which represent opportunities for future work.
5.3. Understanding Opportunities for Improvement.
| Report Length | Citations | ||||
| Chars | Words | Sentences | # Unique Refs | # Inline Citations | |
| Human-written Exemplars | |||||
| Human-written Exemplars | 4381 | 497 | 28 | 23 | 27 |
| Open Source Research Systems | |||||
| DeepResearcher (Llama-4) | 2573 | 319 | 35 | 8 | 7 |
| STORM (Llama-4) | 2766 | 381 | 31 | 18 | 21 |
| OpenScholar (Llama-4) | 3513 | 483 | 26 | 9 | 19 |
| Search AI’s | |||||
| Search AI (Llama-4-Scout) | 1968 | 258 | 16 | 9 | 5 |
| Search AI (GPT-4.1) | 3168 | 404 | 16 | 10 | 61 |
| Search AI (o3) | 3844 | 501 | 24 | 11 | 16 |
| Search AI (Claude) | 3977 | 499 | 27 | 13 | 8 |
| Search AI (Gemini) | 2810 | 395 | 19 | 6 | 8 |
| Commercial Systems | |||||
| OpenAI DeepResearch | 6577 | 864 | 74 | 17 | 6 |
| DeepScholar Baseline | |||||
| DeepScholar-base (Llama-4) | 3864 | 402 | 58 | 21 | 19 |
| DeepScholar-base (GPT-4.1) | 14470 | 1492 | 167 | 19 | 56 |
| DeepScholar-base (GPT-4.1, o3) | 5905 | 642 | 69 | 16 | 20 |
| DeepScholar-base (GPT-4.1, Claude) | 12287 | 1332 | 136 | 23 | 35 |
| DeepScholar-base (GPT-4.1, Gemini) | 6118 | 663 | 81 | 22 | 28 |
| Avg. value over human-written exemplars | Relevant Metric | |
| # Important References from ArXiv.org | 11.47 | Ref. Cov. |
| Median number of citations per reference from ArXiv.org | 647.5 | Doc. Imp. |
In order to further analyze performance and existing opportunities for improvement, we conduct an ablation study, where we consider different retrievers as well as two oracle retriever settings. Table 3 shows these results for two evaluated DeepScholar-base methods, DeepScholar-base (GPT-4.1, Claude) and DeepScholar-base (Llama-4). The table show the performance of either using three different retrieval APIs, including arxiv.org, the default used in our main results, parallel.ai and tavily.com. In addition the table shows to oracle settings for either DeepScholar-base method: the Oracle Retrieval (arxiv.org) setting, provides the system with the ArXiv references from the human-written exemplars labeled as ”Important” following our methodology for evaluating Reference Coverage. The Oracle Retrieval (All) setting, provides the system with any from the human-written exemplars labeled as ”Important” following the same methodology, including both references form ArXiv and once that are found elsewhere.
Overall,the results shown in Table 3 demonstrate that the performance limitations of existing systems for generative research synthesis lie in both their retrieval capabilities to find and select high-quality references sets, as well as their synthesis abilities to surface key facts and extract insights given the retrieved documents.
First, we see that given either oracle retrieval setting, the DeepScholar-base (GPT-4.1, Claude) method nearly saturates performance on Retrieval Quality and Verifiability metrics, whereas the same method using the arxiv.org, parallel.ai or taviliy.com retrievers obtain lower scores on each of these metrics. This finding demonstrates a significant opportunity to improve performance of generative research systems through improvements to the retrieval method. Specifically, existing systems struggle to find a diverse and holistic set of notable sources, reflected by their especially low Reference Coverage and Document Importance scores.
Additionally, we also see that oracle retrieval settings for either DeepScholar-base method attain higher Nugget Coverage, improving the score by up to compared to the rspective arxiv.org, parallel.ai or tavily.ai retreival settings. However, we note that the oracle retreival methods still far from saturate Nugget Coverage, with the DeepScholar-base (GPT-4.1, Claude) Oracle Retrieval (All) attaining a modest score of .. This demonstrates that even with high retrieval quality, existing LLM systems still struggle to effectively surface important facts and synthesize important insights.
5.4. Human Agreement Study
Finally, we study how well our LLM-based evaluation aligns with human judgments, a critical question to validate effectiveness of our fully automated evaluation approach. Overall, we find that each of the metrics we introduce for the DeepScholar-bench task exhibit high agreement between LLM-based judgments and human annotations. We collect over 200 expert annotations, and Table 4 shows the agreement score between human and LLM labelers for organization pairwise comparisons, nugget importance labels, graded relevance scores and reference importance labels. Overall, the results demonstrate above agreement scores across each, reflecting the effectiveness of the DeepScholar-bench evaluation approach.
6. Related Work
Long-form Synthesis Benchmarks. Several recent benchmarks offer datasets designed to evaluate long-form research synthesis tasks, however, their design often involves manually-curated queries, which are prone to data staleness, contamination and limited scalability. DeepScholar-bench addresses these challenges by providing an automated data pipeline and live benchmark with realistic, challenging and recent research synthesis tasks, in contrast to prior works. Specifically, ScholarQABench (Asai et al., 2024) creates realistic literature review questions with detailed answer rubrics, written by expert PhD annotators from the computer science, biomedicine, physics and neuroscience domains. Similarly, OpenResearcher (Zheng et al., 2024) constructs a dataset of around 38 scientific text summarization queries, among other research-style questions, by recruiting experts to write and annotate queries. Likewise, DeepConsult (noa, [n.d.]k) provides a suite of expert-curated, open-ended research queries related to business and consulting. Unfortunately, these expert-curated benchmarks, are expensive to construct, and difficult to update, causing them to quickly become outdated, as new information becomes available, These prior benchmark datasets also risk data contamination, as new models are trained on publically available data. Our work instead proposes a scalable, automated pipeline to provide a live, evergreen dataset that reflects diverse and recent research queries.
Similarly to our approach, the FreshWiki dataset (Shao et al., 2024) is constructed using an automated data pipeline; however, the dataset task focuses on the generation of Wikipedia-like articles, whereas this work focuses on a significantly different and difficult synthesis task based on cutting-edge research. The FreshWiki framework focuses on evaluating the article pre-writing stage as well as the generated full-length article based on a ground truth Wikipedia article and established criteria of a good Wikipedia article (noa, 2025d). This task reflects an interesting area of study; however, in this work, we instead focus on generative research synthesis, and our dataset task focuses on a complex research synthesis task derived from high-quality academic papers. Moreover, our automated evaluation approach is tailored to holistically assess the three key capabilities of generative research synthesis: retrieval, knowledge synthesis and verifiability.
AcademicEval (Zhang et al., 2024) evaluates long-context generation tasks using an ArXiv-derived dataset, similar to ours, however, its task differs substantially, focusing on summarization without retrieval, which is a key component of our task and of generative research synthesis systems more broadly. While the AcademicEval task provides a fixed set of references to a summarization system, our task requires a system to navigate the live web to collect information, and we evaluate this crucial capability of generative research synthesis.
Question Answering Benchmarks. Several recent works on generative research synthesis focus their evaluation on question answering (QA) benchmarks, which, unlike DeepScholar-bench, do not evaluate complex long-form research synthesis tasks and instead focus on short-form answers, whcih can be easily evaluated for correctness. These question answering benchmarks include SimpleQA (Wei et al., 2024), FRAMES (Krishna et al., 2025), GAIA (Mialon et al., 2023), as BrowserComp (Wei et al., 2025), WebWalkerQA (Wu et al., 2025) or others traditionally used to evaluate retrieval-augmented generation (Wadden et al., 2020; Jin et al., 2019; Yang et al., 2018; Joshi et al., 2017; Kwiatkowski et al., 2019; Ho et al., 2020; Trivedi et al., 2022; Lee et al., 2023). While these benchmarks involve a retrieval component, their synthesis task differs substantially from the generative research synthesis task we study. Specifically, each of these QA benchmarks focus on short-form questions with easily verifiable answers and straightforward notions correctness. In contrast, our benchmark provides a framework for studying complex, long-form research synthesis tasks, which lack an absolute notion of correctness and admit many possible reasonable answers.
7. Conclusion
In this work, we introduced DeepScholar-bench, a live dataset and holistic, automated evaluation framework designed to rigorously benchmark an emerging class of systems designed for generative research synthesis. By automatically sourcing queries from high-quality, recent ArXiv papers, our benchmark mitigates the risks of data staleness and training contamination, while offering a real research synthesis task. Moreover, DeepScholar-bench provides an automated evaluation to holistically measure three critical dimensions: retrieval quality, knowledge synthesis and verifiability. We further release DeepScholar-base, a reference pipeline, which we find provides a strong baseline for generative research synthesis. Overall our systematic evalaution of prior open-source systems, search AI’s, OpenAI’s DeepResearch and DeepScholar-base demonstrates significant opportunities for future work, with no system exceeding a score of across all metrics. These results demonstrate both the difficulty of DeepScholar-bench and the exciting opportunity for further advancement in this space. We hope that DeepScholar-bench and DeepScholar-base will support the development of more capable AI systems for generative research synthesis.
Acknowledgements.
This research was supported in part by affiliate members and other supporters of the Stanford DAWN project, including Meta, Google, and VMware, as well as Cisco, SAP, and a Sloan Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.References
- (1)
- noa ([n.d.]a) [n.d.]a. arXiv.org e-Print archive. https://arxiv.org/
- noa ([n.d.]b) [n.d.]b. Claude takes research to new places. https://www.anthropic.com/news/research
- noa ([n.d.]c) [n.d.]c. Gemini Deep Research — your personal research assistant. https://gemini.google/overview/deep-research/
- noa ([n.d.]d) [n.d.]d. Gemini models | Gemini API. https://ai.google.dev/gemini-api/docs/models
- noa ([n.d.]e) [n.d.]e. Grok 3 Beta — The Age of Reasoning Agents | xAI. https://x.ai/news/grok-3
- noa ([n.d.]f) [n.d.]f. Introducing Claude 4. https://www.anthropic.com/news/claude-4
- noa ([n.d.]g) [n.d.]g. Introducing deep research | OpenAI. https://openai.com/index/introducing-deep-research/
- noa ([n.d.]h) [n.d.]h. Introducing Perplexity Deep Research. https://www.perplexity.ai/hub/blog/introducing-perplexity-deep-research
- noa ([n.d.]i) [n.d.]i. Model - OpenAI API. https://platform.openai.com
- noa ([n.d.]j) [n.d.]j. OpenAlex: The open catalog to the global research system | OpenAlex. https://openalex.org/
- noa ([n.d.]k) [n.d.]k. Su-Sea/ydc-deep-research-evals: you.com’s framework for evaluating deep research systems. https://github.com/Su-Sea/ydc-deep-research-evals
- noa ([n.d.]l) [n.d.]l. TREC CAR ← TREC CAR. https://trec-car.cs.unh.edu/
- noa (2025a) 2025a. Llama 4 - a meta-llama Collection. https://huggingface.co/collections/meta-llama/llama-4-67f0c30d9fe03840bc9d0164
- noa (2025b) 2025b. lotus-data/lotus. https://github.com/lotus-data/lotus original-date: 2024-07-16T16:39:06Z.
- noa (2025c) 2025c. meta-llama/Llama-4-Scout-17B-16E-Instruct · Hugging Face. https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E-Instruct
- noa (2025d) 2025d. Wikipedia:Good article criteria. https://en.wikipedia.org/w/index.php?title=Wikipedia:Good_article_criteria&oldid=1303221295 Page Version ID: 1303221295.
- Alzubi et al. (2025) Salaheddin Alzubi, Creston Brooks, Purva Chiniya, Edoardo Contente, Chiara von Gerlach, Lucas Irwin, Yihan Jiang, Arda Kaz, Windsor Nguyen, Sewoong Oh, Himanshu Tyagi, and Pramod Viswanath. 2025. Open Deep Search: Democratizing Search with Open-source Reasoning Agents. https://arxiv.org/abs/2503.20201v1
- Arabzadeh and Clarke (2025) Negar Arabzadeh and Charles L. A. Clarke. 2025. Benchmarking LLM-based Relevance Judgment Methods. In Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (Padua, Italy) (SIGIR ’25). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3726302.3730305
- Asai et al. (2024) Akari Asai, Jacqueline He, Rulin Shao, Weijia Shi, Amanpreet Singh, Joseph Chee Chang, Kyle Lo, Luca Soldaini, Sergey Feldman, Mike D’arcy, David Wadden, Matt Latzke, Minyang Tian, Pan Ji, Shengyan Liu, Hao Tong, Bohao Wu, Yanyu Xiong, Luke Zettlemoyer, Graham Neubig, Dan Weld, Doug Downey, Wen-tau Yih, Pang Wei Koh, and Hannaneh Hajishirzi. 2024. OpenScholar: Synthesizing Scientific Literature with Retrieval-augmented LMs. https://doi.org/10.48550/arXiv.2411.14199 arXiv:2411.14199 [cs].
- Faggioli et al. (2023) Guglielmo Faggioli, Laura Dietz, Charles Clarke, Gianluca Demartini, Matthias Hagen, Claudia Hauff, Noriko Kando, Evangelos Kanoulas, Martin Potthast, Benno Stein, and Henning Wachsmuth. 2023. Perspectives on Large Language Models for Relevance Judgment. In Proceedings of the 2023 ACM SIGIR International Conference on Theory of Information Retrieval. 39–50. https://doi.org/10.1145/3578337.3605136 arXiv:2304.09161 [cs].
- Gao et al. (2023) Tianyu Gao, Howard Yen, Jiatong Yu, and Danqi Chen. 2023. Enabling Large Language Models to Generate Text with Citations. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 6465–6488. https://doi.org/10.18653/v1/2023.emnlp-main.398
- Ho et al. (2020) Xanh Ho, Anh-Khoa Duong Nguyen, Saku Sugawara, and Akiko Aizawa. 2020. Constructing A Multi-hop QA Dataset for Comprehensive Evaluation of Reasoning Steps. In Proceedings of the 28th International Conference on Computational Linguistics, Donia Scott, Nuria Bel, and Chengqing Zong (Eds.). International Committee on Computational Linguistics, Barcelona, Spain (Online), 6609–6625. https://doi.org/10.18653/v1/2020.coling-main.580
- Jin et al. (2019) Qiao Jin, Bhuwan Dhingra, Zhengping Liu, William Cohen, and Xinghua Lu. 2019. PubMedQA: A Dataset for Biomedical Research Question Answering. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 2567–2577. https://doi.org/10.18653/v1/D19-1259
- Joshi et al. (2017) Mandar Joshi, Eunsol Choi, Daniel Weld, and Luke Zettlemoyer. 2017. TriviaQA: A Large Scale Distantly Supervised Challenge Dataset for Reading Comprehension. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Regina Barzilay and Min-Yen Kan (Eds.). Association for Computational Linguistics, Vancouver, Canada, 1601–1611. https://doi.org/10.18653/v1/P17-1147
- Krishna et al. (2025) Satyapriya Krishna, Kalpesh Krishna, Anhad Mohananey, Steven Schwarcz, Adam Stambler, Shyam Upadhyay, and Manaal Faruqui. 2025. Fact, Fetch, and Reason: A Unified Evaluation of Retrieval-Augmented Generation. https://doi.org/10.48550/arXiv.2409.12941 arXiv:2409.12941 [cs].
- Kwiatkowski et al. (2019) Tom Kwiatkowski, Jennimaria Palomaki, Olivia Redfield, Michael Collins, Ankur Parikh, Chris Alberti, Danielle Epstein, Illia Polosukhin, Jacob Devlin, Kenton Lee, Kristina Toutanova, Llion Jones, Matthew Kelcey, Ming-Wei Chang, Andrew M. Dai, Jakob Uszkoreit, Quoc Le, and Slav Petrov. 2019. Natural Questions: A Benchmark for Question Answering Research. Transactions of the Association for Computational Linguistics 7 (2019), 452–466. https://doi.org/10.1162/tacl_a_00276 Place: Cambridge, MA Publisher: MIT Press.
- Kwon et al. (2023) Woosuk Kwon, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody Hao Yu, Joseph E. Gonzalez, Hao Zhang, and Ion Stoica. 2023. Efficient Memory Management for Large Language Model Serving with PagedAttention. https://doi.org/10.48550/arXiv.2309.06180 arXiv:2309.06180 [cs].
- Lee et al. (2023) Yoonjoo Lee, Kyungjae Lee, Sunghyun Park, Dasol Hwang, Jaehyeon Kim, Hong-In Lee, and Moontae Lee. 2023. QASA: Advanced Question Answering on Scientific Articles. In Proceedings of the 40th International Conference on Machine Learning. PMLR, 19036–19052. https://proceedings.mlr.press/v202/lee23n.html
- Li et al. (2025) Dawei Li, Bohan Jiang, Liangjie Huang, Alimohammad Beigi, Chengshuai Zhao, Zhen Tan, Amrita Bhattacharjee, Yuxuan Jiang, Canyu Chen, Tianhao Wu, Kai Shu, Lu Cheng, and Huan Liu. 2025. From Generation to Judgment: Opportunities and Challenges of LLM-as-a-judge. https://doi.org/10.48550/arXiv.2411.16594 arXiv:2411.16594 [cs].
- Li et al. (2024) Ruosen Li, Teerth Patel, and Xinya Du. 2024. PRD: Peer Rank and Discussion Improve Large Language Model based Evaluations. https://doi.org/10.48550/arXiv.2307.02762 arXiv:2307.02762 [cs].
- Liu et al. (2023) Nelson Liu, Tianyi Zhang, and Percy Liang. 2023. Evaluating Verifiability in Generative Search Engines. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 7001–7025. https://doi.org/10.18653/v1/2023.findings-emnlp.467
- Mialon et al. (2023) Grégoire Mialon, Clémentine Fourrier, Craig Swift, Thomas Wolf, Yann LeCun, and Thomas Scialom. 2023. GAIA: a benchmark for General AI Assistants. https://doi.org/10.48550/arXiv.2311.12983 arXiv:2311.12983 [cs].
- Nanni et al. (2017) Federico Nanni, Bhaskar Mitra, Matt Magnusson, and Laura Dietz. 2017. Benchmark for Complex Answer Retrieval. https://doi.org/10.48550/arXiv.1705.04803 arXiv:1705.04803 [cs].
- Patel et al. (2025) Liana Patel, Siddharth Jha, Melissa Pan, Harshit Gupta, Parth Asawa, Carlos Guestrin, and Matei Zaharia. 2025. Semantic Operators: A Declarative Model for Rich, AI-based Data Processing. https://doi.org/10.48550/arXiv.2407.11418 arXiv:2407.11418 [cs].
- Pradeep et al. (2025) Ronak Pradeep, Nandan Thakur, Shivani Upadhyay, Daniel Campos, Nick Craswell, and Jimmy Lin. 2025. The Great Nugget Recall: Automating Fact Extraction and RAG Evaluation with Large Language Models. https://doi.org/10.48550/arXiv.2504.15068 arXiv:2504.15068 [cs].
- Qwen et al. (2025) Qwen, An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, Huan Lin, Jian Yang, Jianhong Tu, Jianwei Zhang, Jianxin Yang, Jiaxi Yang, Jingren Zhou, Junyang Lin, Kai Dang, Keming Lu, Keqin Bao, Kexin Yang, Le Yu, Mei Li, Mingfeng Xue, Pei Zhang, Qin Zhu, Rui Men, Runji Lin, Tianhao Li, Tianyi Tang, Tingyu Xia, Xingzhang Ren, Xuancheng Ren, Yang Fan, Yang Su, Yichang Zhang, Yu Wan, Yuqiong Liu, Zeyu Cui, Zhenru Zhang, and Zihan Qiu. 2025. Qwen2.5 Technical Report. https://doi.org/10.48550/arXiv.2412.15115 arXiv:2412.15115 [cs].
- Rahmani et al. (2024a) Hossein A. Rahmani, Clemencia Siro, Mohammad Aliannejadi, Nick Craswell, Charles L. A. Clarke, Guglielmo Faggioli, Bhaskar Mitra, Paul Thomas, and Emine Yilmaz. 2024a. LLM4Eval: Large Language Model for Evaluation in IR. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval. ACM, Washington DC USA, 3040–3043. https://doi.org/10.1145/3626772.3657992
- Rahmani et al. (2024b) Hossein A. Rahmani, Emine Yilmaz, Nick Craswell, Bhaskar Mitra, Paul Thomas, Charles L. A. Clarke, Mohammad Aliannejadi, Clemencia Siro, and Guglielmo Faggioli. 2024b. LLMJudge: LLMs for Relevance Judgments. https://doi.org/10.48550/arXiv.2408.08896 arXiv:2408.08896 [cs].
- Shao et al. (2024) Yijia Shao, Yucheng Jiang, Theodore A. Kanell, Peter Xu, Omar Khattab, and Monica S. Lam. 2024. Assisting in Writing Wikipedia-like Articles From Scratch with Large Language Models. https://doi.org/10.48550/arXiv.2402.14207 arXiv:2402.14207 [cs].
- Soldaini et al. (2024) Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, Valentin Hofmann, Ananya Jha, Sachin Kumar, Li Lucy, Xinxi Lyu, Nathan Lambert, Ian Magnusson, Jacob Morrison, Niklas Muennighoff, Aakanksha Naik, Crystal Nam, Matthew Peters, Abhilasha Ravichander, Kyle Richardson, Zejiang Shen, Emma Strubell, Nishant Subramani, Oyvind Tafjord, Evan Walsh, Luke Zettlemoyer, Noah Smith, Hannaneh Hajishirzi, Iz Beltagy, Dirk Groeneveld, Jesse Dodge, and Kyle Lo. 2024. Dolma: an Open Corpus of Three Trillion Tokens for Language Model Pretraining Research. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 15725–15788. https://doi.org/10.18653/v1/2024.acl-long.840
- Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. BEIR: A Heterogenous Benchmark for Zero-shot Evaluation of Information Retrieval Models. https://doi.org/10.48550/arXiv.2104.08663 arXiv:2104.08663 [cs].
- Thomas et al. (2024) Paul Thomas, Seth Spielman, Nick Craswell, and Bhaskar Mitra. 2024. Large language models can accurately predict searcher preferences. https://doi.org/10.48550/arXiv.2309.10621 arXiv:2309.10621 [cs].
- Trivedi et al. (2022) Harsh Trivedi, Niranjan Balasubramanian, Tushar Khot, and Ashish Sabharwal. 2022. MuSiQue: Multihop Questions via Single-hop Question Composition. https://doi.org/10.48550/arXiv.2108.00573 arXiv:2108.00573 [cs].
- Upadhyay et al. (2024a) Shivani Upadhyay, Ronak Pradeep, Nandan Thakur, Daniel Campos, Nick Craswell, Ian Soboroff, Hoa Trang Dang, and Jimmy Lin. 2024a. A Large-Scale Study of Relevance Assessments with Large Language Models: An Initial Look. https://doi.org/10.48550/arXiv.2411.08275 arXiv:2411.08275 [cs].
- Upadhyay et al. (2024b) Shivani Upadhyay, Ronak Pradeep, Nandan Thakur, Nick Craswell, and Jimmy Lin. 2024b. UMBRELA: UMbrela is the (Open-Source Reproduction of the) Bing RELevance Assessor. https://doi.org/10.48550/arXiv.2406.06519 arXiv:2406.06519 [cs].
- Voorhees (2009) Ellen M. Voorhees. 2009. I Come Not To Bury Cranfield, but to Praise It. NIST (Oct. 2009). https://www.nist.gov/publications/i-come-not-bury-cranfield-praise-it Last Modified: 2017-02-19T20:02-05:00 Publisher: Ellen M. Voorhees.
- Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or Fiction: Verifying Scientific Claims. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 7534–7550. https://doi.org/10.18653/v1/2020.emnlp-main.609
- Wei et al. (2024) Jason Wei, Nguyen Karina, Hyung Won Chung, Yunxin Joy Jiao, Spencer Papay, Amelia Glaese, John Schulman, and William Fedus. 2024. Measuring short-form factuality in large language models. https://doi.org/10.48550/arXiv.2411.04368 arXiv:2411.04368 [cs].
- Wei et al. (2025) Jason Wei, Zhiqing Sun, Spencer Papay, Scott McKinney, Jeffrey Han, Isa Fulford, Hyung Won Chung, Alex Tachard Passos, William Fedus, and Amelia Glaese. 2025. BrowseComp: A Simple Yet Challenging Benchmark for Browsing Agents. https://doi.org/10.48550/arXiv.2504.12516 arXiv:2504.12516 [cs].
- Worledge et al. (2024) Theodora Worledge, Tatsunori Hashimoto, and Carlos Guestrin. 2024. The Extractive-Abstractive Spectrum: Uncovering Verifiability Trade-offs in LLM Generations. https://doi.org/10.48550/arXiv.2411.17375 arXiv:2411.17375 [cs].
- Wu et al. (2025) Jialong Wu, Wenbiao Yin, Yong Jiang, Zhenglin Wang, Zekun Xi, Runnan Fang, Linhai Zhang, Yulan He, Deyu Zhou, Pengjun Xie, and Fei Huang. 2025. WebWalker: Benchmarking LLMs in Web Traversal. https://doi.org/10.48550/arXiv.2501.07572 arXiv:2501.07572 [cs].
- Yang et al. (2018) Zhilin Yang, Peng Qi, Saizheng Zhang, Yoshua Bengio, William W. Cohen, Ruslan Salakhutdinov, and Christopher D. Manning. 2018. HotpotQA: A Dataset for Diverse, Explainable Multi-hop Question Answering. https://doi.org/10.48550/arXiv.1809.09600 arXiv:1809.09600 [cs].
- Zhang et al. (2024) Haozhen Zhang, Tao Feng, Pengrui Han, and Jiaxuan You. 2024. AcademicEval: Live Long-Context LLM Benchmark. (Oct. 2024). https://openreview.net/forum?id=iRYExPKnxm
- Zheng et al. (2025) Yuxiang Zheng, Dayuan Fu, Xiangkun Hu, Xiaojie Cai, Lyumanshan Ye, Pengrui Lu, and Pengfei Liu. 2025. DeepResearcher: Scaling Deep Research via Reinforcement Learning in Real-world Environments. https://arxiv.org/abs/2504.03160v4
- Zheng et al. (2024) Yuxiang Zheng, Shichao Sun, Lin Qiu, Dongyu Ru, Cheng Jiayang, Xuefeng Li, Jifan Lin, Binjie Wang, Yun Luo, Renjie Pan, Yang Xu, Qingkai Min, Zizhao Zhang, Yiwen Wang, Wenjie Li, and Pengfei Liu. 2024. OpenResearcher: Unleashing AI for Accelerated Scientific Research. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Delia Irazu Hernandez Farias, Tom Hope, and Manling Li (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 209–218. https://doi.org/10.18653/v1/2024.emnlp-demo.22
8. Appendix
8.1. DeepScholar-bench Dataset
We provide a detailed overview and schema of the DeepScholar-bench dataset in Figure 4.
8.2. Baselines
We provide an extended description of each benchmarked method, including relevant implementation details, and parameters used in our evaluation.
8.2.1. DeepResearcher
The DeepResearcher pipeline follows a structured tool-augmented reasoning framework designed for iterative web-based information retrieval. The system mandates explicit reasoning before any tool invocation, with reasoning encapsulated in <think> tags to ensure interpretability and control. After reasoning, the model generates a JSON-formatted request specifying the “web search” tool and its query. These queries are executed via the Lotus Search API, which we replaced with an ArXiv-specific search interface to provide a controlled retrieval API for our evaluation. Retrieved results are returned in a structured format containing the title, URL, and snippet, and are stored in memory for reference across subsequent reasoning steps. This iterative process continues until the model determines that sufficient evidence has been gathered, after which a synthesized final response is produced.
For our experiments, we used Llama-4-Scout-17B-16E-Instruct as the base model, replacing the originally proposed DeepResearcher-7b, since it demonstrated consistently better retrieval-augmented reasoning performance in our experiments. The prompt was slightly modified to align with LLama-4 prompt style as detailed in Box 8.2.1. The retrieval depth was set to 10 sources per query, which is the default in the system and provides a balanced trade-off between coverage and efficiency. We restricted each query to a single rollout with a maximum of 10 steps, following the DeepResearcher defaults; this limit is generous as most rollouts converge in fewer than three steps, but it ensures the system has headroom for more complex queries. The default web search API was replaced with ArXiv search to comply with our benchmark settings.
8.2.2. Openscholar
The OpenScholar pipeline follows a four-stage process: initial retrieval, response and feedback generation, iterative refinement, and citation verification. In the first stage, text segments are retrieved from a fixed index using a contriever model, which encodes texts and retrieves passages based on semantic similarity. These passages are reranked and used to generate an initial draft response, where citations are aligned with the supporting passages. The second stage introduces feedback generation, where the model produces up to three feedback statements highlighting potential improvements in the draft, such as missing content or organization issues; if additional evidence is required, retrieval queries are issued. The third stage iteratively refines the response by conditioning on the previous draft, retrieved passages, and newly added evidence, yielding improved responses at each step until feedback has been fully incorporated. Finally, citation verification ensures that all citation-worthy statements are adequately grounded in the retrieved sources, inserting additional citations where necessary without removing content.
For consistency with other baselines, we employ the Llama-4-Scout-17B-16E-Instruct model for generation. The retrieval pipeline initially collects 100 text segments from peS2o_v3 using the default pes2o_contriever model. The reranker used is OpenScholar_Reranker, also kept at its default setting. To align parameterization across baselines, we increase the number of sources used in generation (top_n) from 10 to 30. Furthermore, the default search API is replaced with the arXiv API, to provided a controlled retreival corpus and API in our experiments.
8.2.3. Search AI
Search AIs are implemented using the open-source OpenDeepSearch (ODS) framework that enables deep web search and retrieval. In particular, the ODS ReAct Agent (instantiated from the smolagents.ToolCallingAgent) is used along with the search agent as an external tool. At each reasoning step, the ReAct agent can either invoke the search agent through a web_search action or decide to produce a final_answer. The search agent interfaces with the search API to fetch relevant academic articles given a query, after which an LLM generates concise summaries of the retrieved content. To maintain consistency with the benchmark setting, the standard search API was replaced with the arXiv API. The regular search agent fails when tasked with full abstract queries; hence the ReAct-based agent was employed, which generates shorter, more effective searchable queries. The agent keeps track of retrieved results across turns, allowing references to past evidence during the reasoning process. After a maximum of 5 iterations, the agent is compelled to conclude with a final response, ensuring bounded computational steps.
For the parameterization of the Search AIs, we set the search agent to retrieve 30 results per query, which is more generous than the default in order to establish fair comparability with other baselines. The reranker parameter was left at infinity, aligning with its default configuration, to avoid prematurely constraining candidate results. The maximum iteration limit was fixed at 5, consistent with the default setup of the ODS framework, providing sufficient exploration without excessive search depth. The ReAct prompt was slightly modified to tailor to the specific use of the ArXiv search API, as presented in Box 8.2.3, 8.2.3 and 8.2.3.
8.2.4. STORM
The STORM pipeline follows a structured multi-stage process to generate comprehensive, Wikipedia-style articles from a given topic. First, related Wikipedia articles are retrieved and their TOCs are clustered to identify candidate perspectives, which act as anchors for exploration. This is followed by Simulated Multi-turn Conversations where an LLM plays both the question-asking and answering roles, querying a retrieval module and synthesizing evidence-based responses. Parallel to this, the model generates a draft outline purely from its parametric knowledge in the Draft Outline Generation stage. The outline is then refined by grounding it with retrieved evidence and conversation outputs. In the final step, each section is drafted with explicit inline citations drawing on both parametric knowledge and retrieved references. All the sections are concatenated together to form the final result.
For parameter settings, we used STORM’s default configurations wherever possible to preserve fidelity to its design: a maximum of 3 turns per perspective, 3 perspectives, and up to 3 search queries per turn. For search, we considered the top 15 results for each query, ensuring a reasonable breadth without overwhelming the pipeline. To make STORM comparable with other baselines, we raised the number of collected references per section title to 30 (more generous than the default), as this allows for richer evidence integration during drafting. Importantly, we replaced the original search API with arXiv search to control the retrieval API for our benchmark settings. Finally, we use Llama-4-Scout-17B-16E-Instruct as the base model.
8.2.5. OpenAI’s DeepResearch
We use OpenAI’s DeepResearch system based on the o3-deep-research model with a custom MCP to only search ArXiv and return results per query. To prevent the model from getting search results after the given paper was uploaded, the MCP used a custom endpoint to set the latest date that it should retrieve. All other settings were set to default values.
8.3. DeepScholar-Base Details and Configurations
DeepScholar-base operates through three main stages: retrieval, filtering, and final generation (Figure 3).
- Retrieval:
In this stage, an LLM generates search queries conditioned on the input abstract and summaries of prior retrievals. Each query is submitted to the configured search API (ArXiv, tavily, etc.) to obtain up to relevant papers within the specified date range. The code and prompt used for this step are provided in Figure 5 and Box 8.3 respectively. This process is repeated times.
1from lotus import web_search23class Query(BaseModel):4 queries: list[str]56# Generate the Queries7queries = get_completion(8 lm,9 query_generation_instruction.format(number_of_queries=num_queries),10 f"Topic: {topic}, Background: {background}",11 response_format=Query,12).queries1314# Search. corpus = ArXiv/Tavily etc.15paper_dfs = []16for query in queries:17 paper_dfs.append(web_search(corpus, query, search_K))1819papers = pd.concat(paper_dfs)
Figure 5. Retrieval stage: query generation and batched search. - Filtering:
Retrieved results are refined using two semantic operators from LOTUS (Patel et al., 2025; noa, 2025b): Sem-Filter and Sem-TopK, which together select the top most relevant papers. The code is given in Figure 6.
1instruction = (2 "given the article’s abstract: {snippet}, "3 "is the article relevant to the specific interests in the user’s query: {user_query}."4)56res_df = docs_df.sem_filter(7 instruction.format(user_query=topic, snippet="{snippet}"),8 strategy="cot"9)1011res_df = res_df.sem_topk(12 instruction.format(user_query=topic, snippet="{snippet}"),13 strategy="cot", k=K14)
Figure 6. Sem-Filter and Sem-TopK code for Filtering Step in DeepScholar-base - Final Generation:
The filtered set of papers is then aggregated via a Sem-Agg query to produce the final output. The corresponding code for this step is shown in Figure 7 with prompt in Box 8.3.
1agg_instruction = section_writer_instructions.format(2 topic=topic,3 section_instructions=section_instructions,4 existing_content=existing_content,5 context="{context}",6)78res: pd.DataFrame = res_df.sem_agg(9 agg_instruction, suffix="summary", group_by=group_by10)
Figure 7. Sem-Agg for final generation in DeepScholar-base
Unless otherwise specified, the pipeline parameters are set to , , , and .
8.4. Evaluation Details
8.4.1. Reference Coverage
Figure 8 illustrates the distribution of important citations across the human exemplar reports in DeepScholar-Bench. For each exemplar, we used the LOTUS program shown in Figure 9 to identify which citations are important and therefore essential to include in a high-quality related work section. We then separate these important citations into two groups: those that appear on ArXiv (shown in red) and those that do not (shown in orange). The blue portion of each bar corresponds to non-essential citations, as determined by the same Lotus-based procedure.
The plot highlights two consistent trends. First, many exemplar related work sections contain a large number of non-essential citations. While such references may be useful for narrative flow or broader context, they are not indispensable. Non-essential citations can be somewhat subjective, depending on how authors choose to frame the story of their paper. In contrast, the important citations represent the “must-have” references i.e., the foundational works in the field that are necessary for situating the contribution. Second, we observe that the red segments (important ArXiv citations) are well distributed across exemplars, indicating that ArXiv is a reliable and sufficiently broad source for recovering many of the essential references.
8.4.2. Ablation Study on Verifiability
In the main paper (Section 3.3, we reported verifiability metrics results using a sliding window of size when computing claim coverage. That is, for each claim sentence, we considered a citation to be valid if any of the references in the same sentence or within one sentence before or after sufficiently supported the claim. Here, we extend this analysis to study the effect of varying the window size. Specifically, we report the citation coverage achieved by different systems when the window size ranges from (same-sentence only) up to (five sentences before or after).
As shown in Figure 10, increasing the window size consistently improves citation coverage across all baselines. This is expected: the larger the window, the higher the probability that one of the cited references in the neighborhood of a claim provides sufficient support. However, we also note that very large window sizes are less desirable in practice, as they often correspond to references being far from the claims they are intended to support, reducing readability and making it harder for readers to verify the connection between claims and citations. Moreover, from Table 5, we see that real academic writing tends to be densely cited, with at least one citation on average per sentence in the human exemplars. Overall, the results of our ablation study highlight the trade-off between stricter precision () and more lenient recall-oriented settings ().
8.4.3. Document Importance Across Human Exemplars
In this section, we illustrate the distribution of document importance, measured by the number of citations of references in the human-written exemplars in DeepScholar-Bench. Figure 11 reports two histograms: (a) the distribution of citation counts across all references, and (b) the distribution restricted to references that appear on ArXiv. We plot the logarithm of citation counts, with values obtained from the OpenAlex API (noa, [n.d.]j), an open and widely used scholarly database that provides citation-level metadata. While citation counts in OpenAlex may not exactly match those from other sources such as Google Scholar, the relative counts are consistent, making it a reliable open-source alternative.
As shown in Figure 11, the distribution is highly skewed due to a small number of papers with exceptionally large citation counts (e.g., over 10k citations). These outliers inflate the mean citation values, resulting in relatively high averages compared to typical references (478.3 citations across all references and 647.6 for ArXiv-only references). In contrast, the median values are lower (31 for all references and 36 for ArXiv-only). This skew highlights the challenge of using citation counts as a proxy for importance, as the median citation count of references, among different human-written exemplars exhibits high variance.